please help
6+6x=7x-4
6x+7x=6-4
13x=2
x=6.5
is this correct
dj








=

Answer by tjnw79(57)

6+6x=7x-4
What you do to one side of the "=" sign you must do the same to the other.
6+6x=7x-4 first add 4 to both sides

(+4)6+6x=7x-4(+4) to get
10+6x=7x now subtract the 6x on both sides
(-6x)10+6x=7x(-6x) now is should look like
10=1x or 10=x